(Related to Checkpoint​ 14.4)  ​(Flotation costs and NPV​ analysis) The Faraway Moving Company is involved in a major plant expansion that involves the expenditure of ​$200 million in the coming year. The firm plans on financing the expansion through the retention of ​$150 million in firm earnings and by borrowing the remaining ​$50 million. In return for helping sell the ​$50 million in new​ debt, the​ firm's investment banker charges a fee of 200 basis points​ (where one basis point is 0.01​ percent). If Faraway decides to adjust for these flotation costs by adding them to the initial​ outlay, what will the initial outlay for the project​ be?

The flotation cost adjusted initial outlay is ​$_______. ​(Round to the nearest​ dollar.)
Answer #1

Investment Banks Fee is =200*0.01 i.e. 2% of $50 million = $1 million

So, initial outlay= (150+50+1) i.e. $201 million.